OpenStudy (anonymous):

Solve for x: x+2/x-2=3/2 (x plus 2 over x minus 2 equals 3 over 2.) A. 2 B. 3 C. 8 D. 10

ganeshie8 (ganeshie8):

As a start : cross multiply

ganeshie8 (ganeshie8):

\[\large \dfrac{x+2}{x-2} =\dfrac{3}{2}\] cross multiplying you get : \[\large 2(x+2) = 3(x-2)\] distribute and isolate x

OpenStudy (anonymous):

so i would cross multiply 2 & 3 & 2&2?

ganeshie8 (ganeshie8):

see this pic http://www.mi2f.com/m/algebra/images/cross-multiply.gif

ganeshie8 (ganeshie8):

i did cross multiplication step already you just need to distribute both sides and isolate x

ganeshie8 (ganeshie8):

\[\large \large 2(x+2) = 3(x-2)\] \[\large \large 2x+4 = 3x-6\]

ganeshie8 (ganeshie8):

see if you can take it home from here